# Simulating Human-like Daily Activities with Desire-driven Autonomy

### v1

@@ Abstract
Existing task-oriented AI agents often depend on explicit instructions or external rewards, limiting their ability to be driven by intrinsic motivations like humans. In this paper, we present a desire-driven autonomy framework to guide a Large Language Model based (LLM-based) agent to simulate human-like daily activities. In contrast to previous agents, our Desire-driven Autonomous Agent (D2A) operates on the principle of intrinsic desire, allowing it to propose and select tasks that fulfill its motivational framework autonomously. Inspired by the Theory of Needs, the motivational framework incorporates an understanding of human-like desires, such as the need for social interaction, personal fulfillment, and self-care. Utilizing a desire-driven task generation mechanism, the agent evaluates its current state and takes a sequence of activities aligned with its intrinsic motivations. Through simulations, we demonstrate that our Desire-driven Autonomous Agent (D2A) generates coherent, contextually relevant daily activities while exhibiting variability and adaptability similar to human behavior. A comparative analysis with other LLM-based frameworks demonstrates that our approach significantly enhances the rationality of the simulated activities  [^1] Project page: https://sites.google.com/view/desire-driven-autonomy.
@@ 1 Introduction
The quest to create human-like autonomous agents (Wang et al., 2016; 2024b) has long been a prominent theme in artificial intelligence (AI) (Fetzer & Fetzer, 1990). A human-like autonomous intelligence is not simply a task executor; it can generate and execute tasks within a motivational framework aligned with its values and goals. Early pioneers such as Alan Turing laid the groundwork for evaluating intelligence through frameworks such as the Turing Test (Pinar Saygin et al., 2000), which posited that a key measure of an agent’s intelligence is its ability to produce behavior indistinguishable from that of a human. Consistent with this perspective, we believe that generating controllable and reasonable human-like behavior is crucial to intelligence. Such capabilities can also enhance user engagement and trust in various applications of AI agents, from service-oriented intelligent assistants that foster realistic interactions (Vrba et al., 2014) to virtual agents in games that create immersive experiences (Hoogendoorn & Soumokil, 2010). Ultimately, the ability to generate human-like activity strengthens the functionality and legitimacy of AI agents in multiple contexts.
@@
The development of large language models (LLMs) (Chang et al., 2024; Achiam et al., 2023) has generated considerable interest in LLM-based agents (Xi et al., 2023; Wang et al., 2024b), demonstrating their significant potential for achieving human-like intelligence. These models benefit from extensive training on diverse datasets, enabling them to possess a wide range of internal knowledge and reasoning abilities. Techniques such as task decomposition, reflection, and memory retrieval can be integrated to further improve agent execution success rates (Shen et al., 2024; Zhang et al., 2024). As a result, LLM-based agents can perform well in diverse contexts and make informed decisions even without specific domain training. However, previous works focus primarily on fixed environments and measure agent performance on predetermined tasks without addressing activity generation (Song et al., 2023). While some approaches use external rewards in open environments to induce more human-like behavior (Sivamayil et al., 2023), the rewards are highly task-specific and fail to model intrinsic motivations that drive more adaptable behaviors, resulting in poor transferability and generalization performance. Research on autonomous agents with diverse, adaptive activity-generation capabilities remains limited.
@@
Our work aims to fill this gap by introducing a desire-driven autonomy framework that empowers agents to generate and pursue activities aligned with their intrinsic motivations, thus enhancing their functionality and relevance in real-world applications. Our idea is inspired by Theory of Needs, which suggest that human behavior can be understood as driven by the motivation to fulfill various intrinsic desires. Consequently, natural human-like behaviors can be produced if an autonomous agent is motivated by a similar system of desires and generates activities accordingly to fulfill them. Specifically, we introduce a framework encompassing dimensions of desires such as hunger’, health’, social connectivity’, and spiritual satisfaction’, aiming to capture a diverse and representative range of human motivations. Instead of instructing the agent to solve clearly defined, closed-ended tasks, we describe desires in these dimensions and ask the agent to try to satisfy these desires. The agent then proposes a series of diverse potential activities and evaluates their impact on various desire dimensions to select the one most aligned with the overall desire. After implementing our framework with the Concordia library (Vezhnevets et al., 2023), we conducted simulations in various textual envionments, from indoor to outdoor setting, from single-agent to multi-agent, with a Game Master providing relevant environment observations to the agent. The experimental results show that our Desire-driven Autonomous Agent (D2A) can generate appropriate activities that effectively and efficiently achieve the described desires. We compared the daily activity sequences generated by D2A with three baseline approaches: ReAct, BabyAGI, and LLMob, which operate based on goal reasoning, agent characteristics, and a combination of both, respectively. As evaluated by GPT-4o and human annotators, D2A generates more human-like daily activities than the baselines according to three criteria: naturalness, coherence, and plausibility.
@@
1. Simulator: We developed a flexible text-based activity simulator based on Large Language Models using Concordia components. The simulator supports various agent types and textual environments, providing reliable interaction and evaluation. As such, it is well-suited for further explorations in human-like activity generation.
@@ 2 Related Work
LLM-based Agents. LLM-based agents (Wang et al., 2024b) have become a thriving area of research due to the impressive problem-solving capabilities of large language models in diverse domains. These agents leverage the extensive training data and internal reasoning capabilities of LLMs to perform tasks in dynamic environments with minimal task-specific fine-tuning. Many existing works focus on goal-driven behaviors, where the agents are designed to decompose user-defined tasks into simpler subtasks and solve them sequentially. For example, systems such as BabyAGI (Nakajima, 2023), AutoGPT (Significant Gravitas, 2023), and Voyager (Wang et al., 2023a) create recursive frameworks for agents to autonomously decompose tasks and execute them in a structured, stepwise manner. These task-oriented agents prioritize the efficient achievement of user-defined goals and rely on external rewards or instructions to guide their actions. However, task-oriented approaches often overlook the importance of intrinsic motivation, a critical aspect of human-like behavior. Some research, such as that by Liu et al. (2023), incorporates long-term memory to improve task coherence across interactions but remains goal-oriented. Other efforts, such as Langchain Agents (Chase, 2022) and CAMEL (Li et al., 2023), enhance the ability of agents to follow instructions and manage task decomposition but do not focus on simulating behaviors that arise from human-like desires or needs. Our research departs from these approaches by focusing on desire-driven autonomy. While most LLM-based agents depend on predefined goals, our framework allows agents to autonomously generate tasks based on a motivational system inspired by Theory of Needs. This shift from external reward-driven behavior to intrinsic desire-based behavior enables agents to simulate more natural, coherent, and human-like daily activities. By aligning agent actions with intrinsic motivations rather than task-specific instructions, we provide a new paradigm for LLM-based agents that fosters more flexible and adaptive behavior in diverse real-world scenarios.
@@ 3 Problem Formulation
In this work, we focus on the problem of simulating human-like daily activities in indoor environments without explicit task instructions. The environment description is denoted by $e$, and the agent profile $p$ includes details such as name, age, and relevant characteristics of the simulated human. For different types of agents, the simulator provides additional customized information $I$ to specify their objectives. For instance, instruction-following agents receive a general goal instruction $g$, while desire-driven agents are assigned desire values $d$. The simulation process consists of $T$ time steps. At each time step $t$, the agent must generate a new activity $a_{t}$ based on all the given context, including past activities $a_{0:t-1}$ and observations $o_{0:t-1}$:
@@
4 Human-like Daily Activities Simulator
Our simulator is built on Concordia (Vezhnevets et al., 2023), a text-based simulator empowered by large language models. The Concordia library equips our simulator with a rich set of components, each serving specific roles such as goal and plan management, time reporting, memory storage, retrieval, and more. It also provides wrapped interfaces that support various open-source models and APIs, allowing the simulator to switch between different underlying models. These components form the core of Concordia’s interaction framework, enabling flexible communication between the agent and the Game Master, which is the environment controller.
Building on this agent-controller framework, we introduce several enhancements designed for human-like daily activity simulation. First, environment consistency is enforced through explicit procedures and prompts, ensuring agents only interact with items present in the environment, thus maintaining realism. This consistency enables the dynamic addition or removal of interactable items within text-based environments, allowing precise control over the agent’s activity space. Second, our simulator supports self-reflection during both action execution and state updates, improving decision-making and reliability. Finally, we provide robust methods for evaluation and visualization, enabling the assessment of both single-step actions and entire activity sequences for coherence and plausibility. These custom modifications facilitate human-like daily activity simulations, ensuring adaptability, stability, and reliability across a wide range of agent types and behaviors.
5 Desire-driven Autonomy for Daily Activity Simulation
Our Desire-driven autonomy framework consists of two core modules: the Value System and the Desire-driven Planner, supported by essential components such as memory and text-based environments. The Value System manages desire components, each representing the level of satisfaction for a predefined desire dimension. During the simulation setup, these desire components are initialized with initial values $v_{d}^{0}\in\mathbb{R}$ and expected values $v_{d}^{*}\in\mathbb{R}$, where $d$ represents a specific desire dimension. In each simulation step $t$, the Desire-driven Autonomous Agent (D2A) performs four key procedures using the Value System and the Desire-driven Planner to generate the activity $a_{t}$: Qualitative Value Description, Activity Proposal, Activity Evaluation, and Activity Selection. At the end of each step, the desire components update $v_{d}^{t}$ based on $v_{d}^{t-1}$, the executed activity $a_{t}$, and the observation $o_{t}$ provided by the Game Master. The goal of the agent is to perform appropriate activities so its final desire components $v_{d}^{T}$ can match the expected values $v_{d}^{*}$.
@@ 5.1 Desire Generation
The theory of needs suggests that individual behavior as being motivated by meeting the individual’s needs or wants, such as safety, social, esteem, and self-actualization. Furthermore, individuals may prioritize different needs based on their unique profiles, traits, and personalities. These considerations underscore the complexity of human needs and illustrate how these needs drive behavior. To simulate human behavior, we developed a value system that maps personality profile $p$ with the expected desire values $v^{*}$ (See Appendix E and F). Specifically, for the indoor, single-agent environment, we define 11 desire dimensions. The physiological dimension, representing fundamental survival needs, is further divided into categories such as ‘hunger’, ‘thirst’, ‘sleepiness’, ‘cleanliness’, ‘comfort’, and ‘health’. Safety needs are encapsulated by the ‘safeness’ dimension. To capture love and belonging needs, which reflect the desire for relationships, we introduce the ‘social connectivity’ dimension. Finally, the self-actualization needs, linked to personal growth, are categorized into ‘joy’, ‘passion’, and ‘spiritual satisfaction’. For the outdoor party multi-agents environment, we further introduce ‘recognition’, ‘sense of control’ and ‘sense of superiority’ three desire dimensions to direct activity generation with higher-level motivations in social environment. To quantitatively track changes in the level of these desires and qualitatively translate numerical values into descriptive states, we apply a [0-10] Likert scale for each dimension, mapping numerical values into textual descriptive outcomes (See Appendix G).
@@
5.2 Value System
As shown in Figure 2, the Value System executes three key procedures at each simulation step: Quantitative Value Deduction, Qualitative Value Description, and Value Update. The Quantitative Value Deduction procedure models the fluctuating nature of human desires by incorporating a decay mechanism for all intrinsic values. At the beginning of each time step $t$, the agent’s intrinsic value $v_{d}^{t-1}$ for desire dimension $d$ is reduced by 1 with a probability $p_{d}$. This probability is determined by the agent’s profile characteristics, following the same mechanism used for the expected desire value mapping described in Appendix E and F. Since fulfilling these desires serves as D2A’s intrinsic motivation, the Quantitative Value Deduction procedure is designed to drive the agent to continuously generate desire-driven activities.
Since many studies have shown that large language models often struggle with interpreting numerical values, we designed the Qualitative Value Description procedure to translate these values into meaningful textual desire descriptions. As shown in our ablation studies (Sec. 6.3.3), the descriptions are important for the agent to understand its current desire states so as to generate appropriate activities. To achieve this, we implemented a workflow that converts all intrinsic numerical values $v^{t}$ into descriptive desire sentences $D^{t}$ by asking a set of predefined questions (Appendix H) to backbone LLMs at each simulation step $t$. The agent is expected to recognize the shortfall $\delta^{t}$ between its current desires and the expected values, based on the descriptive desire $D^{t}$ and its internal profile $p$ throughout the simulation.
@@ 5.3 Desire-Driven Planner
Inspired by the work of Tree of Thoughts (Yao et al., 2024), we argue that the paradigm of presenting multiple candidate activities and evaluating them helps the agent select the best action. Thus, during the Activity Proposal procedure (Appendix K), we prompt the agent to generate $N$ candidates (also referred to as planner width, which is set to 3 in our default experimental configuration) activities that may have positive impacts on its desires given $p$, $e$, $a_{0:t-1}$, $o_{0:t-1}$, and desire descriptions $D^{t}$:
@@
(Here and below, we use conditional probability-like form to represent the sampling formula. This is not a strict probabilistic sampling as in mathematical derivations, but rather a formal way to illustrate the information provided as context and the type of content that needs to be sampled.)
These $N$ candidate activities are then passed on to the Activity Evaluation procedure (Appendix L). The Planner evaluates each activity $a_{t}^{i}$ by imagining the resulting desire states $\hat{D_{t}^{i}}$ that the agent would experience after taking action $a_{t}^{i}$:
@@
Finally, in the Activity Selection procedure (Appendix M), the agent compares the forecasted desire states $\hat{D_{t}^{i}}$ for all $N$ proposed actions. Based on these evaluations, the Planner selects the activity $a_{t}$ that is expected to best fulfill the agent’s intrinsic desires:
@@
This process ensures that the chosen activity optimally aligns with the agent’s intrinsic motivations. After the activity is selected and executed, the environment produces the corresponding observation $o_{t}$, and the Value System updates the agent’s desire components to reflect the new state, completing the simulation step.

### v2

@@ Abstract
Desires motivate humans to interact autonomously with the complex world. In contrast, current AI agents require explicit task specifications, such as instructions or reward functions, which constrain their autonomy and behavioral diversity. In this paper, we introduce a Desire-driven Autonomous Agent (D2A) that can enable a large language model (LLM) to autonomously propose and select tasks, motivated by satisfying its multi-dimensional desires. Specifically, the motivational framework of D2A is mainly constructed by a dynamic Value System, inspired by the Theory of Needs. It incorporates an understanding of human-like desires, such as the need for social interaction, personal fulfillment, and self-care. At each step, the agent evaluates the value of its current state, proposes a set of candidate activities, and selects the one that best aligns with its intrinsic motivations. We conduct experiments on Concordia, a text-based simulator, to demonstrate that our agent generates coherent, contextually relevant daily activities while exhibiting variability and adaptability similar to human behavior. A comparative analysis with other LLM-based agents demonstrates that our approach significantly enhances the rationality of the simulated activities  [^1] Project page: https://sites.google.com/view/desire-driven-autonomy.
@@ 1 Introduction
The quest to create human-like autonomous agents (Wang et al., 2016; 2024b) has long been a prominent theme in artificial intelligence (AI) (Fetzer, 1990). A human-like autonomous intelligence is not simply a task executor; it can generate and execute tasks within a motivational framework aligned with its values and goals. Early pioneers such as Alan Turing laid the groundwork for evaluating intelligence through frameworks such as the Turing Test (Pinar Saygin et al., 2000), which posited that a key measure of an agent’s intelligence is the ability to produce behavior indistinguishable from that of a human. Consistent with this perspective, we believe that generating controllable and reasonable human-like behavior is crucial to agents. Such capabilities can also enhance user engagement and trust in various applications of AI agents, from service-oriented intelligent assistants that foster realistic interactions (Vrba et al., 2014) to virtual agents in games that create immersive experiences (Hoogendoorn & Soumokil, 2010). Ultimately, the ability to generate human-like activity strengthens the functionality and legitimacy of AI agents in multiple contexts.
@@
The remarkable progress in large language models (LLMs) (Chang et al., 2024; Achiam et al., 2023) has spurred considerable research into LLM-based agents (Xi et al., 2025; Wang et al., 2024b), revealing their potential for achieving human-like intelligence. Extensive training on diverse datasets equips LLMs with a broad range of internal knowledge and reasoning abilities. Techniques such as task decomposition, reflection, and memory retrieval can be integrated to further improve agents’ task execution success rates (Shen et al., 2024; Zhang et al., 2024). As a result, LLM-based agents can perform well in diverse contexts and make informed decisions even without specific domain training. However, previous works focus primarily on fixed environments and measure agent performance on predetermined tasks without addressing activity generation (Song et al., 2023). While some approaches use external rewards in open environments to induce more human-like behavior (Sivamayil et al., 2023), the rewards are highly task-specific and fail to model intrinsic motivations that drive more adaptable behaviors, resulting in poor transferability and generalization ability. Research on autonomous agents with diverse, adaptive activity-generation capabilities remains limited.
@@
Our work aims to fill this gap by introducing a desire-driven autonomy framework that empowers agents to generate and pursue activities aligned with their intrinsic motivations, thus enhancing their functionality in real-world applications. Our idea is inspired by the Theory of Needs (Maslow, 1943; McClelland, 1987), which suggests that human behavior can be understood as driven by the motivation to fulfill various intrinsic desires. Consequently, natural human-like behaviors can be produced if an autonomous agent is motivated by a similar system of desires and generates activities accordingly to fulfill them. Specifically, we introduce a framework that includes dimensions of desire such as hunger, health, social connectivity, and spiritual satisfaction, aiming to encompass a broad and representative spectrum of human motivations. Instead of instructing the agent to solve clearly defined tasks, we describe the desires in these dimensions and ask the agent to attempt to satisfy these desires. The agent then proposes a set of candidate activities and evaluates their impact on various desire dimensions to select the one most aligned with the overall desire.
We implement our Desire-driven Autonomous Agent (D2A) across diverse textual environments, simulating activities in both indoor and outdoor settings, as well as in single-agent and multi-agent scenarios. The environments are based on Concordia (Vezhnevets et al., 2023), with a Game Master providing relevant observations to the agent. The experimental results show that our agents can generate appropriate activities that effectively and efficiently satisfy the described desires. We compared the daily activity sequences generated by D2A with three baseline agents: ReAct, BabyAGI, and LLMob, which operate based on goal reasoning, agent characteristics, and a combination of both, respectively. As evaluated by GPT-4o and human annotators, D2A generates more human-like daily activities than the baselines according to naturalness, coherence, and plausibility.
@@
1. Text-based Daily Activity Simulator: We developed a flexible text-based activity simulator based on Large Language Models using Concordia components. The simulator supports various agent types and textual environments, providing reliable interaction and evaluation. As such, it is well-suited for further explorations in human-like activity generation.
@@ 2 Related Work
LLM-based Agents. LLM-based agents (Wang et al., 2024b) have become a thriving area of research because of the impressive problem-solving capabilities of large language models in diverse domains. These agents leverage the extensive training data and internal reasoning capabilities of LLMs to perform tasks in dynamic environments with minimal task-specific fine-tuning. Many existing works focus on goal-driven behaviors, where the agents are designed to decompose user-defined tasks into simpler subtasks and solve them sequentially. For example, systems such as BabyAGI (Nakajima, 2023), AutoGPT (Significant Gravitas, 2023), and Voyager (Wang et al., 2023a) create recursive frameworks for agents to autonomously decompose tasks and execute them in a structured, stepwise manner. These task-oriented agents prioritize the efficient achievement of user-defined goals and rely on external rewards or instructions to guide their actions. However, task-oriented approaches often overlook the importance of intrinsic motivation, a critical aspect of human-like behavior. Some research, such as that by Liu et al. (2023), incorporates long-term memory to improve task coherence across interactions but remains goal-oriented. Other efforts, such as Langchain Agents (Chase, 2022) and CAMEL (Li et al., 2023), enhance the ability of agents to follow instructions and manage task decomposition but do not focus on simulating behaviors that arise from human-like desires or needs. Our research departs from these approaches by focusing on desire-driven autonomy. While most LLM-based agents depend on predefined goals, our framework allows agents to autonomously generate tasks based on a motivational system inspired by the Theory of Needs. This shift from external reward-driven behavior to intrinsic desire-based behavior enables agents to simulate more natural, coherent, and human-like daily activities. By aligning agent actions with intrinsic motivations rather than task-specific instructions, we provide a new paradigm for LLM-based agents that fosters more flexible and adaptive behavior in diverse real-world scenarios.
@@ 3 Problem Formulation
In this work, we focus on the problem of simulating human-like daily activities in interactive environments without explicit task instructions. The environment description is denoted by $e$, and the agent profile $p$ includes details such as name, age, and relevant characteristics of the simulated human. For different types of agents, the simulator provides additional customized information $I$ to specify their objectives. For instance, instruction-following agents receive a general goal instruction $g$, while desire-driven agents are assigned desire values $d$. The simulation process consists of $T$ time steps. At each time step $t$, the agent must generate a new activity $a_{t}$ based on all the given context, including past activities $a_{0:t-1}$ and observations $o_{0:t-1}$:
@@
4 Text-based Daily Activity Simulator
Our simulator is built on Concordia (Vezhnevets et al., 2023), a text-based simulator empowered by large language models. The Concordia library equips our simulator with a rich set of components, each serving specific roles such as goal and plan management, time reporting, memory storage, retrieval, and more. It also provides wrapped interfaces that support various open-source models and APIs, allowing the simulator to switch between different underlying models. These components form the core of Concordia’s interaction framework, enabling flexible communication between the agent and the environment controller (Game Master).
Building on this framework, we introduce two main enhancements for daily activity simulation. First, environment consistency is enforced through explicit procedures and prompts, ensuring agents only interact with items present in the environment, thus maintaining realism. This consistency enables the dynamic addition or removal of items within text-based environments, allowing precise control over the agent’s activity space. Second, we provide robust methods for evaluation and visualization, enabling the assessment of both single-step actions and entire activity sequences for coherence and plausibility. These custom modifications facilitate the daily activity simulations, ensuring adaptability, stability, and reliability across a wide range of agent types and behaviors.
5 Desire-driven Autonomous Agent
Our Desire-driven autonomy framework consists of two core modules: the Value System and the Desire-driven Planner, supported by other essential components for agents, such as memory and self-reflection. The Value System manages desire components, each representing the level of satisfaction for a predefined desire dimension. During the simulation setup, these desire components are initialized with initial values $v_{d}^{0}\in\mathbb{R}$ and expected values $v_{d}^{*}\in\mathbb{R}$, where $d$ represents a specific desire dimension. In each simulation step $t$, the Desire-driven Autonomous Agent (D2A) performs four key procedures using the Value System and the Desire-driven Planner to generate the activity $a_{t}$: Qualitative Value Description, Activity Proposal, Activity Evaluation, and Activity Selection. At the end of each step, the values update $v_{d}^{t}$ based on $v_{d}^{t-1}$, the executed activity $a_{t}$, and the observation $o_{t}$ provided by the Game Master. The goal of the agent is to perform appropriate activities so its final desires $v_{d}^{T}$ can match the expected values $v_{d}^{*}$.
@@ 5.1 Desire Generation
The theory of needs suggests that individual behavior is motivated by meeting the individual’s needs or wants, such as safety, social, esteem, and self-actualization. Furthermore, individuals may prioritize different needs based on their unique profiles, traits, and personalities. These considerations underscore the complexity of human needs and illustrate how these needs drive behavior. To simulate human behavior, we developed a value system that maps personality profile $p$ with the expected desire values $v^{*}$ (See Appendix E and F). Specifically, for the indoor, single-agent environment, we define 11 dimensions of desired value. The physiological dimension, representing fundamental survival needs, is further divided into categories such as ‘hunger’, ‘thirst’, ‘sleepiness’, ‘cleanliness’, ‘comfort’, and ‘health’. Safety needs are encapsulated by the ‘safeness’ dimension. To capture love and belonging needs, which reflect the desire for relationships, we introduce the ‘social connectivity’ dimension. Finally, the self-actualization needs, linked to personal growth, are categorized into ‘joy’, ‘passion’, and ‘spiritual satisfaction’. For the outdoor party multi-agents environment, we further introduce ‘recognition’, ‘sense of control’, and ‘sense of superiority’ three desire dimensions to steer activity generation with higher-level motivations in a social environment. To quantitatively track changes in the level of these desires and qualitatively translate numerical values into descriptive states, we apply a [0-10] Likert scale for each dimension, mapping numerical values into textual descriptive outcomes (See Appendix G).
@@
5.2 Value System for Desire Evaluation
As shown in Figure 2, the Value System executes three key procedures at each simulation step: Quantitative Value Deduction, Qualitative Value Description, and Value Update. The Quantitative Value Deduction procedure models the fluctuating nature of human desires by incorporating a decay mechanism for all intrinsic values. At the beginning of each time step $t$, the agent’s intrinsic value $v_{d}^{t-1}$ for desire dimension $d$ is reduced by 1 with a probability $p_{d}$. This probability is determined by the agent’s profile characteristics, following the same mechanism used for the expected desire value mapping described in Appendix E and F. Since fulfilling these desires serves as D2A’s intrinsic motivation, the Quantitative Value Deduction procedure drives the agent to continuously generate desire-driven activities.
Since many studies have shown that large language models often struggle with interpreting numerical values, we designed the Qualitative Value Description procedure to translate these values into meaningful textual desire descriptions. As shown in our ablation studies (Sec. 6.4), the descriptions are important for the agent to understand its current desire states so as to generate appropriate activities. To achieve this, we implemented a workflow that converts all intrinsic numerical values $v^{t}$ into descriptive desire sentences $D^{t}$ by asking a set of predefined questions (Appendix H) to backbone LLMs at each simulation step $t$. The agent is expected to recognize the shortfall $\delta^{t}$ between its current desires and the expected values, based on the descriptive desire $D^{t}$ and its internal profile $p$ throughout the simulation.
@@ 5.3 Desire-Driven Planner
Inspired by the work of Tree of Thoughts (Yao et al., 2024), we argue that the paradigm of presenting multiple candidate activities and evaluating them enables the agent to select the best action. Thus, during the Activity Proposal procedure (Appendix J.2), we prompt the agent to generate $N$ candidates (also referred to as planner width, which is set to 3 in our default experimental configuration) activities that may have positive impacts on its desires given $p$, $e$, $a_{0:t-1}$, $o_{0:t-1}$, and desire descriptions $D^{t}$:
@@
(Here and below, we use a conditional probability-like form to represent the sampling formula. This is not a strict probabilistic sampling as in mathematical derivations, but rather a formal way to illustrate the information provided as context and the type of content that needs to be sampled.)
These $N$ candidate activities are then passed on to the Activity Evaluation procedure (Appendix J.3). The Planner evaluates each activity $a_{t}^{i}$ by imagining the resulting desire states $\hat{D_{t}^{i}}$ that the agent would experience after taking action $a_{t}^{i}$:
@@
Finally, in the Activity Selection procedure (Appendix J.4), the agent compares the forecasted desire states $\hat{D_{t}^{i}}$ for all $N$ proposed actions. Based on these evaluations, the Planner selects the activity $a_{t}$ that is expected to best fulfill the agent’s intrinsic desires:
@@
This process ensures that the chosen activity optimally aligns with the agent’s intrinsic motivations. After the activity is selected and executed, the environment produces the corresponding observation $o_{t}$, and the Value System for desire evaluation updates to reflect the new state, completing the simulation step.